Two numbers 2^59 and 3^x are added ,if 5 is at the unit's place of the result, then value of x is?
Answer by Alan3354(69443)

Two numbers 2^59 and 3^x are added ,if 5 is at the unit's place of the result, then value of x is?
-------------
2^59 has a 1's digit = 8
3^x has to have a 7
-------
x = 3 + 4n, n = 0,1,2,3...